Draw the graphs representing equations 2x=y+2 and 4x+3y=24 on the same graph paper.
Answer by mananth(16949)

2 x -1 y = 2 .............1
4 x + 3 y = 24 .............2
equation1 2 = 2
when x= 0 y= -2 ( 0 , -2 ) -2
when y = 0 x= 1 ( 1 , 0 ) 1
Equation 2
when x= 0 y= 8 ( 0 , 8 ) 8
when y = 0 x= 6 ( 6 , 0 )